When x number of units are sold, the price of each unit (in
dollars) is given by p = -x/2 + 75. Find the unit price when the following quantities
are sold: 2, 7, 9, 11.

Answer by checkley71(8403)

p=-x/2+75
p=-2/2+75
p=-1+75
p=74 answer.
p=-11/2+75
p=-5.5+75
p=69.5 answer.
--------------------------------
I DID 2 TO SHOW YOU HOW TO SOLVE THESE TYPE OF PROBLEMS.
NOW YOU SOLVE THE REMAINING 2 USING THE SAME PROCEDURE.